OpenStudy (anonymous):

GIVING A MEDAL!!!!!!!! if h(x)=(f o g)(x) and h(x)=5(x+1)^3, find f(x) and g(x)

OpenStudy (luigi0210):

Simple.. replace \( x+1\) with \(x\)

OpenStudy (kirbykirby):

\((f \,\,o \,\,g)(x)=f\left(g(x)\right)=h(x)\) So, thing of two function, such then when you plug in "g(x)" into your f(x) function, you get h(x).

OpenStudy (kirbykirby):

think of two functions*

OpenStudy (kirbykirby):

such that* ugh so many typos

OpenStudy (anonymous):

but I don't get what you have to plug in

OpenStudy (kirbykirby):

Ok let's look at a simple example: If you have \(f(x)=x^2+7\), then \(f(5)=5^2+7\), you plug in "x=5" Now, let say \(f(x)=x^2+7\) and \(g(x)=x^5\), then \(f(g(x))=f(x^5)\), so you plug in \(x^5\) where you see \(x\). So, \(f(g(x))=(x^5)^2+7=x^{10}+7\).

OpenStudy (kirbykirby):

In your example, you want to find \(f(x)\) and \(g(x)\) such that \(f(g(x))=5(x+1)^3\)